In the absence of added quenchers, the fluorescence quantum yield is found to be 30%. Calculate the fluorescence lifetime under these conditions on the assumption that the fluorescence decay is monoexponential.

10 mM NaI causes the fluorescence yield and lifetime both to decrease to 10% of their original values. Calculate the Stern-Vollmer quenching constant.Calculate the bimolecular rate constant for the quenching by NaI.
Answer #1

It is known that quantum yield is given by
Φ = Lifetime (τ0) x rate of emission (T)
Since it is given that the fluorescence decay is exponential and Φ is 0.3 (30%)
So, the rate of emission for a common fluorophore is 1x108 s-1
Using this the lifetime of fluorophore is
0.3/1x108 s-1 = 0.3x108 s = 3ns

Quenching of fluorescence by a quencher can be explained by Stern-Volmer equation;

F0/F = 1 + KSV [Q]

Where, F0 and F are the fluorescence intensities in the absence and presence of quencher.

              [Q] is the concentration of quencher.

              KSV is the stern-Volmer quenching constant;

                    KSV = k * τ0

                           Here, k is the bimolecular quenching constant and τ0 is the lifetime of the fluorophore in the absence of quencher.

Also, the fluorescence intensity(F) is proportional to the amount of light absorbed and the fluorescence quantum yield(Φ). It is given by the equation

              F = kIoΦA

                        Where, I0 is the incident light intensity and A is the absorbance.

Therefore, the ratio of fluorescence intensities can be converted into ratio of fluorescence quantum yield.

Φ0/Φ = 1 + KSV [Q]

Given, concentration of quencher = 10mM = 0.01M

            Φ0 = 0.3

             Φ = 0.03

ie, 0.3/0.03 = 1 + KSV * 0.01

Stern-Volmer constant, KSV = 900

And bimolecular quenching constant, k = KSV/ τ0

                                                                            = 900/(3*10-9)

= 30*1010 M-1 s-1